OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x – 3)(x + 1)(x – 5)?

OpenStudy (babyslapmafro):

What values, when substituted for x, produce a value of zero. In other words, what values produce f(x)=0?

OpenStudy (anonymous):

3,-1,5?

OpenStudy (babyslapmafro):

yes

OpenStudy (anonymous):

can you help check a few that ii did ?

OpenStudy (anonymous):

x-3=0, x+1=0, x-5=0 x=3, x=-1, x=5

OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = –2x3 + 5x2 – 6x + 4 ? Positive: 1; Negative: 2 or 0; Complex 2 or 0 What is the quotient: (2x2 + 2x – 24) ÷ (x + 4) ? 2x – 6 What are the zeros of the polynomial function: f(x) = (x – 3)(x + 1)(x – 5)? 3, –1, 5 Using synthetic division, what is the quotient: (x3 + 4x2 – 9x – 8) ÷ (x – 2) ? x2 + 6x + 3 + r(-2) What are the possible rational zeros of f(x) = 3x3 – 4x2 – 7x + 21 ? ± 1, ± 3, ± 7, ± 21 What polynomial has roots of –4, –2, and 5 ? x3 + x2 – 22x – 40

OpenStudy (anonymous):

would someone mind checking those ?